OpenStudy (anonymous):

NEED HELP!!! Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (4,2)

OpenStudy (anonymous):

y-2=2(x-4) simplify it

OpenStudy (anonymous):

Where did you get y-2?

OpenStudy (anonymous):

So far I have 2x-8

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) the point is (x1,y1) and m is the slope

OpenStudy (anonymous):

the one i wrote is the standard form..

OpenStudy (anonymous):

standard form of a line is ax+by=c

OpenStudy (anonymous):

there are many forms of line, dear @insanitycat

OpenStudy (anonymous):

How would I solve y-2=2x-8 ?

OpenStudy (anonymous):

yes but there are names for each slope intercept y=mx+b point-slope : y-y1=m(x-x1) and standard:ax+by=c

OpenStudy (anonymous):

you would add 2 to both sides and then you have your line

OpenStudy (anonymous):

that would be your equation

OpenStudy (anonymous):

Actually all I am looking for is the equation in slope-intercept

OpenStudy (anonymous):

So what would the equation be?

OpenStudy (anonymous):

y=2x-6 <---this is slope intercept form (y=mx+b)

OpenStudy (anonymous):

Thanks that is what I got . Now do I simplify it? to y=x-3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

since for that you would have to divide the entire equation by 2 and you would have y/2=x-3